A bag contains 5 red balls and 6 white balls. If we draw 4 balls, find the probability that at least 3 balls are white.
1
Expert's answer
2020-03-09T10:56:58-0400

1)


"P(W=4)=\frac{\binom{6}{4}}{\binom{11}{4}}"

"\binom{6}{4}=\frac{6!}{4!2!}=15"

"\binom{11}{4}=\frac{11!}{4!7!}=330"

2)


"P(W=3)=\frac{\binom{6}{3}}{\binom{11}{4}}"

"\binom{6}{3}=\frac{6!}{3!3!}=20"

3)


"P(W\ge3)=\frac{15+20}{330}=\frac{7}{66}\approx 0.106"